59mandatory:

Type the correct answer in the box. Use numerals instead of words. What is the value of this expression when a = -5 and b = -1? For a = -5 and b = -1, the expression has a value of

Spectrum:

What's the expression?

59mandatory:

@spectrum -5 and -1

Lovers:

copy and paste the entire problem

Spectrum:

Two variables is not an expression.

Lovers:

i mean the ENTIRE thing

59mandatory:

ohh the other part didnt paste one sec

59mandatory:

(\root(3)(a3))/(b)

Lovers:

screenshot it and post it here

59mandatory:

alr

Lovers:

it might be better that way

59mandatory:

59mandatory:

@spectrum

Lovers:

yeah, i can't help with this, sorry

59mandatory:

its alr

Lovers:

what type of math even is ths ?

59mandatory:

root

Lovers:

no, i mean the class

59mandatory:

algebra 2

Spectrum:

Take the values of your variables and plug them into the equation.

Lovers:

@oliver69

YourlocalRandom2000:

3x + 2y = 3(4) + 2(5) (substitute x = 4 and y = 5) = 12 + 10 = 22 So the value of the expression 3x + 2y when x = 4 and y = 5 is 22.